OpenStudy (anonymous):

What is the slope of the line that passes through the points (16, –1) and (–4, 10)?

OpenStudy (kropot72):

\[Slope\ m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}=\frac{-1-10}{-16-(-4)}=you\ can\ calculate\]